OpenStudy (anonymous):

find the minimum value of\[f(x)=x+\frac{1}{x}+x^2+\frac{1}{x^2}\]

OpenStudy (noelgreco):

How can you find relative extreme values?

OpenStudy (datanewb):

Taking the derivative doesn't seem to help. \[f'(x) = 1 - \frac{1}{x^2} + 2x - \frac{2}{x^3}\] Only thing I can tell is that it's shooting towards infinity on either side of zero, and on the ends as well. Interested in following this thread to see what the answer is. :) The minimum must be close to -1?

OpenStudy (datanewb):

f'(-1) = 0 and \[f''(x) = \frac{2}{x^3}+2+\frac{6}{x^4}\\f''(-1) > 0\], so -1 is at least a local minimum.

OpenStudy (anonymous):

\[t=x+\frac{1}{x}\]\(|t|\ge2\)\[f(t)=t^2+t-2\]finding minimum of this

OpenStudy (datanewb):

Well, the minimum of f(t) is -2, then \[-2 = x + \frac{1}{x} \\ x = -1,f(x) = 0\] very cool! How much more difficult would it be to find the min of say \[f(x)=x+\frac{1}{x}+x^2+\frac{2}{x^2}\]?

OpenStudy (anonymous):

well this one is impossible to solve only with pen and paper :)